# Quasinormal modes of Schwarzschild-de Sitter black holes in semi-open systems

### v1

@@ V Exceptional points
A widely accepted conclusion is that generating a second-order EP (two-fold degeneracy of eigenvalues) typically requires two free real parameters Yang:2025dbn. When there are insufficient free parameters, the phenomenon of EPs generally does not occur, and it is instead replaced by mode repulsion or avoided crossing Berti:2025hly; Dias:2021yju; Motohashi:2024fwt; Oshita:2025ibu; Lo:2025njp; Takahashi:2025uwo. For example, considering the QNM spectra of the Kerr black hole, there is only one dimensionless parameter, i.e., $a/M$. For $0.875\leq a/M\leq 0.915$, the mode $\omega_{225}$ resonates with the mode $\omega_{226}$, and for $0.95\leq a/M\leq 1-10^{-6}$, the mode $\omega_{315}$ resonates with the mode $\omega_{316}$ Motohashi:2024fwt. The physical quantity used to quantitatively characterize resonance phenomena is called the excitation factor (EF), which has been widely studied in many gravitational systems Leaver:1986gd; Sun:1988tz; Andersson:1995zk; Glampedakis:2003dn; Berti:2006wq; Zhang:2013ksa; Silva:2024ffz; Oshita:2024wgt; Oshita:2025ibu; Lo:2025njp; Takahashi:2025uwo; Oshita:2021iyn; Chen:2024hum; Motohashi:2024fwt; Kubota:2025hjk; Wu:2025sbq. As such type of the repulsion phenomenon occurs, the excitation factor will become very large for some appropriate parameters. At the same time, either the real parts of the modes cross and the imaginary parts repel, or the imaginary parts of the modes cross and the real parts repel. Therefore, if only one parameter is varied, the avoided crossing occurs.
@@
However, when the number of free (real) parameters becomes $2$, the story changes. For example, in Cavalcante:2024kmy; Cavalcante:2024swt, a scalar field of mass $\mu$ around a Kerr black hole is considered. For $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$, it is found that the longest-living mode and the first overtone coincide, and this pair of parameters is called the EP. The hysteresis phenomenon of spectra also occurs, which is interpreted through a thermodynamic analogy. One can refer more details to the recent work Cavalcante:2025abr. Returning to our study, in Sec. III, we have obtained the QNM spectra migration due to the reflectivity $\mathcal{K}$ varies in the interval $[0,1]$. Therein, given the position of the wall $x_{0}$, and the cosmological constant via $r_{c}$ with the unit being $r_{e}=1$, the only parameter that can be freely adjusted is $\mathcal{K}\in\mathbb{R}$. So, as expected, there is no EP as $\mathcal{K}$ varies (at least for $\mathcal{K}\in[0,1]$). However, in this section, we allow the reflectivity $\mathcal{K}$ to be a complex constant, $\mathcal{K}\in\mathbb{C}$, which can provide two free real parameters. An imaginary part of the reflectivity denotes that the reflection process around the event horizon imparts a phase change of the wave, as opposed to a pure change in intensity. Doing so would lead to the appearance of exceptional points. This is also the first implementation of EPs caused by changes in QNM boundary conditions in the gravity aspect.
@@
Now, we will search for the EP of $\mathcal{K}$ on the complex plane. The QNM sepctra results for the real $\mathcal{K}>0$ are cornerstones of our EP finding approach. The schematic diagram for the EP finding approach is given in Fig. 6. On the one hand, for a small $|\mathcal{K}|_{\text{min}}$, one starts from the left black point in Fig. 6, in order to implement the outcomes of the QNM spectra on the circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{min}}$. On the other hand, for a large $|\mathcal{K}|_{\text{max}}$, one starts from the right black point in Fig. 6, in order to implement the outcomes of the QNM spectra on the circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{max}}$. Continuously narrow the gap between $|\mathcal{K}|_{\text{min}}$ and $|\mathcal{K}|_{\text{max}}$ so that the mode repulsion phenomenon shown in Fig. 7 occurs, in which we have chosen typical parameters. From the left panel of Fig. 7, as the complex parameter $\mathcal{K}$ move along the circumference $|\mathcal{K}|_{\text{min}}$, the fundamental mode $n=0$ goes back the original position, and the overtone $n=2$ migrates to the position for the original overtone $n=1$. From the right panel of Fig. 7, as the complex parameter $\mathcal{K}$ moves along the circumference $|\mathcal{K}|_{\text{max}}$, the fundamental mode $n=0$ migrates to the position for the original overtone $n=1$, and the overtone $n=2$ migrates to the original position for the fundamental mode $n=0$. One can find a peculiar parameter $\mathcal{K}_{\text{EP}}$ so that the red solid line is closest to the green dashed line regardless of in the left panel or in the right panel of Fig. 7. Such point, which is depicted as a red star, is nothing but the exceptional point. For the parameter in Fig. 7, the EP for the fundamental mode and the second overtone is given by $\mathcal{K}_{\text{EP}}\simeq 2.42704\times 10^{-5}-6.00589\times 10^{-5}\mathrm{i}$. It is worth noting that in the process of moving the parameter, no exchange between two modes $n=0$ and $n=2$ actually occurred. This phenomenon arises because the loops, $|\mathcal{K}|=|\mathcal{K}|_{\text{min}}$ and $|\mathcal{K}|=|\mathcal{K}|_{\text{max}}$, actually contains infinitely many EPs Cavalcante:2025abr, rather than just one. Some of these EPs involves the $n=0$ mode or $n=2$ mode coalescence with another different modes. The method given above can actually be considered a relatively reliable way to find EPs, especially for cases where the equations of QNM spectra can be analytically established.
@@
Figure 7: The migrations for fundamental mode $n=0$ (red solid line) and the overtone $n=2$ (green dashed line) within the parameter $r_{e}=1$, $r_{c}=1.1$, $x_{0}=-150$, where the left panel corresponds to the parameter circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{min}}$ while the right panel corresponds to the parameter circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{max}}$. The red point, blue point and green point represent the positions for the original fundamental mode $n=0$, the original first overtone $n=1$ and the original second overtone $n=2$, respectively.
@@
where $R_{\mathcal{K}}$ is the radius of the small circumference, and $\theta$ is the angle parameter. The parameter range is $0,4\pi]$, which is equivalent to the parameter circling around the EP twice. We take parameters $r_{e}=1$, $r_{c}=1.1$ and $x_{0}=-150$ as an example to illustrate, where one of the EPs is $\mathcal{K}_{\text{EP}}\simeq 2.42704\times 10^{-5}-6.00589\times 10^{-5}\mathrm{i}$, and other parameters are also treated similarly. In left panel of Fig. 8, the migrations of the fundamental modes are shown, in which different colors represent different radii. In fact, the results of the left panel of Fig. 8 is also the migration of the second overtone, where the difference is that the mode $n=2$ starts from $\theta=2\pi$ and ends at $\theta=2\pi$. The left of the Fig. 8 shows the exchange of modes, which is known as the so-called the hysteresis phenomenon of QNMs Cavalcante:2024kmy; Cavalcante:2024swt; Cavalcante:2025abr; Motohashi:2024fwt. Furthermore, regarding black hole spectroscopy, we will demonstrate that the necessity of applying Puiseux series near the EP in the parameterized analysis of QNMs Cao:2025afs. In order to show it, for a given $R_{\mathcal{K}}$, averaging the distance between $\omega_{0}$ and $\omega_{2}$ with $\theta\in[0,4\pi]$, where the dashed double arrow line represents this distance depicted in Fig. 8. The derived average gives the quantity which is denoted as $|\omega_{0}-\omega_{2}|$. It can be plotted that the relation between $|\omega_{0}-\omega_{2}|$ and $(R_{\mathcal{K}})^{1/2}$ in Fig. 8. From the Fig. 8, we fit the numerical results of $|\omega_{0}-\omega_{2}|$ and $R_{\mathcal{K}}$ to see that $|\omega_{0}-\omega_{2}|$ is proportional to $R_{\mathcal{K}}^{1/2}$, giving the structure of the Puiseux series at around the EP Ashida:2020dkc [cf. Eq. (111) therein].
Figure 8: The left panel shows the migrations for fundamental mode $n=0$ with the parameter $r_{e}=1$, $r_{c}=1.1$, $x_{0}=-150$, for the parametric equation (32) of the small circumference. Different color lines correspond to the different radii. On each lines, the symbol $\blacksquare$ corresponds to $\theta=0$ and $\theta=2\pi$, the symbol $\bullet$ corresponds to $\theta=\pi/2$, $\pi$, $3\pi/2$, $5\pi/2$, $3\pi$ and $7\pi/2$ and $\theta=2\pi$. The magenta star is the degenerate QNM spectra for the EP. In addition, the migration for second overtone starts from $\theta=2\pi$. $|\omega_{0}-\omega_{2}|$ v.s. $(R_{\mathcal{K}})^{1/2}\equiv|\mathcal{K}-\mathcal{K}_{\text{EP}}|^{1/2}$ is shown in right panel, and the dashed line is corresponding linear fit.
@@ VI Conclusions and discussion
In conclusion, this work employs a combined analytic (Heun function) and numerical approach to comprehensively elucidate the impact of semi-open systems on the perturbation of SdS black holes. The extension of this work is concentrating on the Kerr-de Sitter black holes in semi-open systems Hatsuda:2020sbn. For this rotation context, exceptional points and lines Cao:2025afs caused by boundary conditions can be studied, and even higher-order exceptional points could be found. The ringdown waveform corresponding to EPs caused by boundary conditions can be further studied, similar to Yang:2025dbn; PanossoMacedo:2025xnf.
@@ Appendix A The Heun’s equation
Table 1: Singular points of Eq. (33), the corresponding coefficients $a_{0},b_{0}$ in the index equation, and the associated indices $(\rho_{1},\rho_{2})$ for scalar $(s=0)$, electromagnetic $(s=1)$, and axial gravitational $(s=2)$ perturbations.
@@
where $\gamma+\delta+\epsilon=\alpha+\beta+1$. One can apply the Frobenius method to get the infinite power series solutions. In particular, two local solutions at $z=0$ are given by Oshita:2022pyf; Hatsuda:2020sbn; Oshita:2021iyn
@@
$q=\frac{r_{c}(-r_{c}-2r_{e})\Big[-3l^{2}-3l-\Lambda r_{c}(r_{c}+r_{e})\Big]}{\Lambda(-r_{c}-r_{e})(r_{c}-r_{e})(-r_{c}^{2}-2r_{c}r_{e})}\,,\quad\alpha=1\,,\quad\beta=1\,.$ (44)
@@
$q=-\frac{3l(l+1)r_{c}(-r_{c}-2r_{e})}{\Lambda(-r_{c}-r_{e})(r_{c}-r_{e})(-r_{c}^{2}-2r_{c}r_{e})}\,,\quad\alpha=0\,,\quad\beta=2\,.$ (45)
@@
$q=-\frac{3r_{c}(-r_{c}-2r_{e})\Big[l^{2}+l-\Lambda r_{c}(r_{c}+r_{e})\Big]}{\Lambda(-r_{c}-r_{e})(r_{c}-r_{e})(-r_{c}^{2}-2r_{c}r_{e})}\,,\quad\alpha=-1\,,\quad\beta=3\,.$ (46)

### v2

@@ V Exceptional points
A widely accepted conclusion is that generating a second-order EP (two-fold degeneracy of eigenvalues) typically requires two free real parameters Yang et al. (2025a). When there are insufficient free parameters, the phenomenon of EPs generally does not occur, and it is instead replaced by mode repulsion or avoided crossing Berti and others (2025); Dias et al. (2022); Motohashi (2025); Oshita et al. (2025); Lo et al. (2025); Takahashi et al. (2025). For example, considering the QNM spectra of the Kerr black hole, there is only one dimensionless parameter, i.e., $a/M$. For $0.875\leq a/M\leq 0.915$, the mode $\omega_{225}$ resonates with the mode $\omega_{226}$, and for $0.95\leq a/M\leq 1-10^{-6}$, the mode $\omega_{315}$ resonates with the mode $\omega_{316}$ Motohashi (2025). The physical quantity used to quantitatively characterize resonance phenomena is called the excitation factor (EF), which has been widely studied in many gravitational systems Leaver (1986); Sun and Price (1988); Andersson (1995); Glampedakis and Andersson (2003); Berti and Cardoso (2006); Zhang et al. ; Silva et al. (2024); Oshita and Cardoso (2025); Oshita et al. (2025); Lo et al. (2025); Takahashi et al. (2025); Oshita (2021); Chen et al. (2025b); Motohashi (2025); Kubota and Motohashi (2025); Wu et al. (2025b). As such type of the repulsion phenomenon occurs, the excitation factor will become very large for some appropriate parameters. At the same time, either the real parts of the modes cross and the imaginary parts repel, or the imaginary parts of the modes cross and the real parts repel. Therefore, if only one parameter is varied, the avoided crossing occurs.
@@
However, when the number of free (real) parameters becomes $2$, the story changes. For example, in Cavalcante et al. (2024b, a), a scalar field of mass $\mu$ around a Kerr black hole is considered. For $(M\mu)_{c}\simeq 0.3704981$ and $(a/M)_{c}\simeq 0.9994660$, it is found that the longest-living mode and the first overtone coincide, and this pair of parameters is called the EP. The hysteresis phenomenon of spectra also occurs, which is interpreted through a thermodynamic analogy. One can refer more details to the recent work Cavalcante et al. (2025). Returning to our study, in Sec. III, we have obtained the QNM spectra migration due to the reflectivity $\mathcal{K}$ varies in the interval $[0,1]$. Therein, given the position of the wall $x_{0}$, and the cosmological constant via $r_{c}$ with the unit being $r_{e}=1$, the only parameter that can be freely adjusted is $\mathcal{K}\in\mathbb{R}$. So, as expected, there is no EP as $\mathcal{K}$ varies (at least for $\mathcal{K}\in[0,1]$). However, in this section, we allow the reflectivity $\mathcal{K}$ to be a complex constant, $\mathcal{K}\in\mathbb{C}$, which can provide two free real parameters. An imaginary part of the reflectivity denotes that the reflection process around the event horizon imparts a phase change of the wave, as opposed to a pure change in intensity. Doing so would lead to the appearance of exceptional points. This is also the first implementation of EPs caused by changes in QNM boundary conditions in the gravity aspect.
@@
Now, we will search for the EP of $\mathcal{K}$ on the complex plane. The QNM sepctra results for the real $\mathcal{K}>0$ are cornerstones of our EP finding approach. The schematic diagram for the EP finding approach is given in Fig. 6. On the one hand, for a small $|\mathcal{K}|_{\text{min}}$, one starts from the left black point in Fig. 6, in order to implement the outcomes of the QNM spectra on the circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{min}}$. On the other hand, for a large $|\mathcal{K}|_{\text{max}}$, one starts from the right black point in Fig. 6, in order to implement the outcomes of the QNM spectra on the circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{max}}$. Continuously narrow the gap between $|\mathcal{K}|_{\text{min}}$ and $|\mathcal{K}|_{\text{max}}$ so that the mode repulsion phenomenon shown in Fig. 7 occurs, in which we have chosen typical parameters. From the left panel of Fig. 7, as the complex parameter $\mathcal{K}$ move along the circumference $|\mathcal{K}|_{\text{min}}$, the fundamental mode $n=0$ goes back the original position, and the overtone $n=2$ migrates to the position for the original overtone $n=1$. From the right panel of Fig. 7, as the complex parameter $\mathcal{K}$ moves along the circumference $|\mathcal{K}|_{\text{max}}$, the fundamental mode $n=0$ migrates to the position for the original overtone $n=1$, and the overtone $n=2$ migrates to the original position for the fundamental mode $n=0$. One can find a peculiar parameter $\mathcal{K}_{\text{EP}}$ so that the red solid line is closest to the green dashed line regardless of in the left panel or in the right panel of Fig. 7. Such point, which is depicted as a red star, is nothing but the exceptional point. For the parameter in Fig. 7, the EP for the fundamental mode and the second overtone is given by $\mathcal{K}_{\text{EP}}\simeq 2.42704\times 10^{-5}-6.00589\times 10^{-5}\mathrm{i}$. It is worth noting that in the process of moving the parameter, no exchange between two modes $n=0$ and $n=2$ actually occurred. This phenomenon arises because the loops, $|\mathcal{K}|=|\mathcal{K}|_{\text{min}}$ and $|\mathcal{K}|=|\mathcal{K}|_{\text{max}}$, actually contains infinitely many EPs Cavalcante et al. (2025), rather than just one. Some of these EPs involves the $n=0$ mode or $n=2$ mode coalescence with another different modes. The method given above can actually be considered a relatively reliable way to find EPs, especially for cases where the equations of QNM spectra can be analytically established.
@@
Figure 7: The migrations for fundamental mode $n=0$ (red solid line) and the overtone $n=2$ (green dashed line) within the parameter $r_{e}=1$, $r_{c}=1.1$, $x_{0}=-150$, where panel (a) corresponds to the parameter circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{min}}$ while panel (b) corresponds to the parameter circumference $|\mathcal{K}|=|\mathcal{K}|_{\text{max}}$. The red point, blue point and green point represent the positions for the original fundamental mode $n=0$, the original first overtone $n=1$ and the original second overtone $n=2$, respectively.
@@
where $R_{\mathcal{K}}$ is the radius of the small circumference, and $\theta$ is the angle parameter. The parameter range is $0,4\pi]$, which is equivalent to the parameter circling around the EP twice. We take parameters $r_{e}=1$, $r_{c}=1.1$ and $x_{0}=-150$ as an example to illustrate, where one of the EPs is $\mathcal{K}_{\text{EP}}\simeq 2.42704\times 10^{-5}-6.00589\times 10^{-5}\mathrm{i}$, and other parameters are also treated similarly. In left panel of Fig. 8, the migrations of the fundamental modes are shown, in which different colors represent different radii. In fact, the results of the left panel of Fig. 8 is also the migration of the second overtone, where the difference is that the mode $n=2$ starts from $\theta=2\pi$ and ends at $\theta=2\pi$. The left of the Fig. 8 shows the exchange of modes, which is known as the so-called the hysteresis phenomenon of QNMs Cavalcante et al. (2024b, a, 2025); Motohashi (2025). Furthermore, regarding black hole spectroscopy, we will demonstrate that the necessity of applying Puiseux series near the EP in the parameterized analysis of QNMs Cao et al. (2025a). In order to show it, for a given $R_{\mathcal{K}}$, averaging the distance between $\omega_{0}$ and $\omega_{2}$ with $\theta\in[0,4\pi]$, where the dashed double arrow line represents this distance depicted in Fig. 8. The derived average gives the quantity which is denoted as $|\omega_{0}-\omega_{2}|$. It can be plotted that the relation between $|\omega_{0}-\omega_{2}|$ and $(R_{\mathcal{K}})^{1/2}$ in Fig. 8. From the Fig. 8, we fit the numerical results of $|\omega_{0}-\omega_{2}|$ and $R_{\mathcal{K}}$ to see that $|\omega_{0}-\omega_{2}|$ is proportional to $R_{\mathcal{K}}^{1/2}$, giving the structure of the Puiseux series at around the EP Ashida et al. (2021) [cf. Eq. (111) therein].
Figure 8: Panel (a) shows the migrations for fundamental mode $n=0$ with the parameter $r_{e}=1$, $r_{c}=1.1$, $x_{0}=-150$, for the parametric equation (32) of the small circumference. Different color lines correspond to the different radii. On each lines, the symbol $\blacksquare$ corresponds to $\theta=0$ and $\theta=2\pi$, the symbol $\bullet$ corresponds to $\theta=\pi/2$, $\pi$, $3\pi/2$, $5\pi/2$, $3\pi$ and $7\pi/2$ and $\theta=2\pi$. The magenta star is the degenerate QNM spectra for the EP. In addition, the migration for second overtone starts from $\theta=2\pi$. $|\omega_{0}-\omega_{2}|$ v.s. $(R_{\mathcal{K}})^{1/2}\equiv|\mathcal{K}-\mathcal{K}_{\text{EP}}|^{1/2}$ is shown in panel (b), and the dashed line is corresponding linear fit.
@@ VI Conclusions and discussion
In conclusion, this work employs a combined analytic (Heun function) and numerical approach to comprehensively elucidate the impact of semi-open systems on the perturbation of SdS black holes. The extension of this work is concentrating on the Kerr-de Sitter black holes in semi-open systems Hatsuda (2020). For this rotation context, exceptional points and lines Cao et al. (2025a) caused by boundary conditions can be studied, and even higher-order exceptional points could be found. The ringdown waveform corresponding to EPs caused by boundary conditions can be further studied, similar to Yang et al. (2025a); Panosso Macedo et al. (2025).
@@ Appendix A The Heun’s equation
Table 1: Singular points of Eq. (33), the corresponding coefficients $a_{0},b_{0}$ in the index equation $\rho(\rho-1)+a_{0}\rho+b_{0}=0$, and the associated indices $(\rho_{1},\rho_{2})$ for scalar $(s=0)$, electromagnetic $(s=1)$, and axial gravitational $(s=2)$ perturbations.
@@
where $\gamma+\delta+\epsilon=\alpha+\beta+1$. One can apply the Frobenius method to get the infinite power series solutions. In particular, two local solutions at $z=0$ are given by Oshita et al. (2022); Hatsuda (2020); Oshita (2021)
@@
$q=\frac{-3\ell^{2}-3\ell-\Lambda r_{c}(r_{c}+r_{e})}{\Lambda(-r_{c}-r_{e})(r_{c}-r_{e})}\,,\quad\alpha=1\,,\quad\beta=1\,.$ (44)
@@
$q=-\frac{3\ell(\ell+1)}{\Lambda(-r_{c}-r_{e})(r_{c}-r_{e})}\,,\quad\alpha=0\,,\quad\beta=2\,.$ (45)
@@
$q=-\frac{3\Big[\ell^{2}+\ell-\Lambda r_{c}(r_{c}+r_{e})\Big]}{\Lambda(-r_{c}-r_{e})(r_{c}-r_{e})}\,,\quad\alpha=-1\,,\quad\beta=3\,.$ (46)
@@
Appendix C Quasinormal modes for $s=0$
Here we show a concrete result of QNMs for $s=0$ to clarify that the QNMs of other spins will have similar qualitative properties. In Fig. 9, we show the migrations of the QNMs in the cases with $s=0$. From this figure, it can be seen that the qualitative migrations of QNMs are similar to the case of $s=2$.
Figure 9: For the scalar perturbation $s=0$, the migrations of the QNM spectra with $\mathcal{K}\in[0,1]$ varying in the complex plane for the first 1010 modes, where the parameters are chosen $r_{e}=1$, $r_{c}=1.1$, and $x_{0}=-100$. Different color lines correspond to different modes.